# Unable to Forget: Proactive Interference Reveals Working Memory Limits in LLMs Beyond Context Length

### v2

@@ Abstract
Information retrieval in Large Language Models (LLMs) is increasingly recognized as intertwined with generation capabilities rather than mere lookup. While longer contexts are often assumed to improve retrieval, the effects of intra-context interference remain understudied. To address this, we adapt the proactive interference (PI) paradigm from cognitive science, where earlier information disrupts recall of newer updates. In humans, susceptibility to such interference is inversely linked to working memory capacity. We introduce PI-LLM, an evaluation that sequentially streams semantically related key–value updates and queries only the final values. Although these final values are clearly positioned just before the query, LLM retrieval accuracy declines log-linearly toward zero as interference accumulates; errors arise from retrieving previously overwritten values. Attempts to mitigate interference via prompt engineering (e.g., instructing models to ignore earlier input) yield limited success. These findings reveal a fundamental constraint on LLMs’ ability to disentangle interference and flexibly manipulate information, suggesting a working memory bottleneck beyond mere context access. This calls for approaches that strengthen models’ ability to suppress irrelevant content during retrieval.
Machine Learning, ICML
@@ 1 Introduction
Figure 1: Universal log-linear decline in retrieval performance due to interference. Increasing the amount of interfering information (i.e., the number of semantically similar items) preceding a retrieval target within a language model’s input context results in a log-linear decrease in retrieval accuracy across diverse models. The target is positioned after the interfering information and explicitly referenced in the prompt to reduce search difficulty and isolate interference effects. Retrieval accuracy declines continuously as models retrieve information located after increasing amounts of interference (x-axis: log-scaled, base 10).
Current research indicates that Large Language Models (LLMs) generally struggle with retrieval tasks when closely related pieces of information are present (Vodrahalli et al., 2024). Furthermore, reasoning models do not effectively improve performance in these scenarios (OpenAI, 2025a). However, most studies —having already labeled retrieval as a ‘long-context’ challenge— prioritize input length as the primary determinant of retrieval difficulty, relegating other factors to a secondary role.
@@
  • Universal Log-Linear Decay Reveals Unified Interference Limit. Across all state-of-the-art models, retrieval accuracy declines log-linearly toward zero as update count, number of tracked keys, or value length increases (Figures 1,8,9). This indicates a robust, negative log-linear relationship between retrieval performance and interference load and points to a unified capacity limit for interference tolerance.
@@ 2.2 Results and Discussion
Although prompt choice shifts absolute accuracy (He et al., 2024), our study focuses on the trend of decline. This emphasis on relative trends, rather than raw scores, neutralizes single-prompt variability. By testing many models from multiple vendors—and repeating the analysis with additional prompt templates—we find the decay slope is strikingly stable (see Figure  11 ’Relevance meta-prompt’).
@@ 5.3 Hack method: Mock-QA-reset succeeds
This limitation is further highlighted by additional experiments with natural-language prompts designed to instruct LLMs to ignore or forget prior information—such as the ’soft session reset’ shown in Figure 10—which were largely ineffective (see performance in Figure  11). These findings indicate that current LLMs cannot be reliably controlled through explicit natural-language user instructions alone; precise, natural-language-based adjustments of memory and attention remain an open challenge.

### v3

@@ Abstract
Information retrieval in Large Language Models (LLMs) is increasingly recognized as intertwined with generation capabilities rather than mere lookup. While longer contexts are often assumed to improve retrieval, the effects of intra-context interference remain understudied. To address this, we adapt the proactive interference (PI) paradigm from cognitive science, where earlier information disrupts recall of newer updates. In humans, susceptibility to such interference is inversely linked to working memory capacity. We introduce PI-LLM, an evaluation that sequentially streams semantically related key–value updates and queries only the final values. Although these final values are clearly positioned just before the query, LLM retrieval accuracy declines log-linearly toward zero as interference accumulates; errors arise from retrieving previously overwritten values. Attempts to mitigate interference via prompt engineering (e.g., instructing models to ignore earlier input) yield limited success. These findings reveal a fundamental constraint on LLMs’ ability to disentangle interference and flexibly manipulate information, suggesting a working memory bottleneck beyond mere context access. This calls for approaches that strengthen models’ ability to suppress irrelevant content during retrieval. Code and data are publicly available at github.com/zhuangziGiantfish/Unable-to-Forget.
@@ 1 Introduction
Figure 1: Universal log-linear decline in retrieval performance due to interference. Increasing the amount of interfering information preceding a retrieval target within a language model’s input context results in a log-linear decrease in retrieval accuracy across diverse models. The target is positioned after the interfering information and explicitly referenced in the prompt to reduce search difficulty and isolate interference effects. (x-axis: the number of semantically similar items, log-scaled; asterisk: MoE models)
Evaluation for information retrieval in Large Language Models (LLMs) primarily uses input length as the main measure of task difficulty. Current research indicates that language models generally struggle with retrieval tasks when closely related pieces of information are present (Vodrahalli et al., 2024). Furthermore, reasoning models do not effectively improve performance in these scenarios (OpenAI, 2025a).
@@
  • Universal Log-Linear Decay Reveals Unified Interference Limit. Across all state-of-the-art models, retrieval accuracy declines log-linearly toward zero as update count, number of tracked keys, or value length increases (Figures 1,9,10). This indicates a robust, negative log-linear relationship between retrieval performance and interference load and points to a unified capacity limit for interference tolerance.
@@ 2.2 Results and Discussion
Robustness to Prompt Variations
Although absolute retrieval accuracy can shift with changes in prompt wording (He et al., 2024), our study emphasizes the relative trend of performance decline rather than raw accuracy scores. This approach effectively neutralizes variability arising from individual prompt formulations.
To further confirm the robustness of the observed proactive interference (PI) effect, we tested additional prompt templates explicitly designed to verify task comprehension. Specifically, we introduced meta-relevant prompts that first ask the LLM to articulate the ”task mission”—for example, explicitly prompting the model to ”describe the goal of this task” before retrieval. This step ensured the models fully understood the retrieval objective of identifying ”the last value” (see Figure 12, ‘Relevance meta-prompt’). Across these prompt variations, the qualitative trend of performance decline—specifically, the log-linear decay in accuracy—remained consistently robust.
Figure 4: Step-like failure pattern in sequential key–value update tests. Retrieval accuracy remains near-perfect as interfering information is added in strictly sequential order, until a model-specific threshold is reached—after which performance drops abruptly to near-zero. Within the same model family, larger models exhibit a higher threshold (better capacity). Despite quantitative differences, all models show the same two-plateau, step-function pattern, reflecting a hard capacity limit. This stands in contrast to the gradual log-linear decay observed under random update order (see Fig. 1). (x-axis: number of interfering items, log-scaled; asterisk: MoE models)
Additionally, we rearranged the input organization to test PI under both randomly shuffled and strictly sequential update sequences (i.e., sequential key–value updates without randomization; see Figure 4). Notably, in sequential mode, retrieval accuracy remains stable until reaching a model-specific interference threshold, after which performance sharply and consistently drops to near-zero—a two-plateau, step-like pattern contrasting with the gradual log-linear decay observed in random mode. The consistent interference-induced decline across diverse models and input structures further underscores the robustness and generalizability of the observed PI phenomenon.
@@ 5.3 Hack method: Mock-QA-reset succeeds
This limitation is further highlighted by additional experiments with natural-language prompts designed to instruct LLMs to ignore or forget prior information—such as the ’soft session reset’ shown in Figure 11—which were largely ineffective (see performance in Figure  12). These findings indicate that current LLMs cannot be reliably controlled through explicit natural-language user instructions alone; precise, natural-language-based adjustments of memory and attention remain an open challenge.
@@ 9.2.1 Connecting Behavioral Evidence with Mechanistic Interpretability
Code Availability
The code for all experiments is publicly available at: https://github.com/zhuangziGiantfish/Unable-to-Forget
@@
Acknowledgments
Special thanks to Sam Zheyang Zheng, Simons Foundation’s Flatiron Institute CCN researcher and NYU Ph.D. candidate, whose clear, contemplative discussions with Chupei(CW)—linking human and LLM focus to dynamical systems—encouraged the initiation of this investigation.
We also thank Yikun Kuang (Ph.D. student, Center for Data Science–CILVR Lab, NYU), whose consultations—from early brainstorming to conference advice—enriched our work by offering a machine learning perspective that complemented our cognitive approaches.